OpenStudy (anonymous):

Marisa painted one-half of her bedroom in three-fourths of an hour. At this rate, how long would it take her the paint the entire room?

OpenStudy (texaschic101):

(1/2) / (3/4) = 1/x cross multiply 1/2x = 3/4 x = (3/4) / (1/2) x = 3/4 * 2 x = 6/4 = 1 1/2 hrs (thats 1 and 1/2 hrs)...or 1.5 hrs